### v2

# Detecting Flaky Tests in Quantum Software: A Dynamic Approach

Abstract
Flaky tests, tests that pass or fail nondeterministically without changes to code or environment, pose a serious threat to software reliability. While classical software engineering has developed a rich body of dynamic and static techniques to study flakiness, corresponding evidence for quantum software remains limited. Prior work relies primarily on static analysis or small sets of manually reported incidents, leaving open questions about the prevalence, characteristics, and detectability of flaky tests.
This paper presents the first large-scale dynamic characterization of flaky tests in quantum software. We executed the Qiskit terra test suite 10,000 times across 23 releases in controlled environments. For each release, we measured test-outcome variability, identified flaky tests, estimated empirical failure probabilities, analyzed recurrence across versions, and used Wilson confidence intervals to quantify rerun budgets for reliable detection. We further mapped flaky tests to terra subcomponents to assess component-level susceptibility.
Across 27,026 test cases, we identified 290 distinct flaky tests. Although overall flakiness rates were low (0–0.4%), flakiness was highly episodic: nearly two-thirds of flaky tests appeared in only one release, while a small subset recurred intermittently or persistently. Many flaky tests failed with very small empirical probabilities ($\hat{p}\approx 10^{-4}$), implying that tens of thousands of executions may be required for confident detection. Flakiness was unevenly distributed across subcomponents, with transpiler and quantum_info accounting for the largest share.
These results show that quantum test flakiness is rare but difficult to detect under typical continuous integration budgets. To support future research, we release a public dataset of per-test execution outcomes.
keywords:
quantum flaky tests , dynamic test execution , quantum software testing , quantum software engineering
[…]
[aff1]organization=University of Maryland, Baltimore County, city=Baltimore, state=MD, country=USA
[…]
[aff2]organization=Toronto Metropolitan University, city=Toronto, state=ON, country=Canada
1 Introduction
Flaky tests, i.e., tests that unpredictably alternate between passing and failing despite no changes to code or environment, pose a persistent threat to software reliability (Luo et al., 2014; Micco, 2017; Memon et al., 2017). Their presence reduces trust in test outcomes, obscures real defects, and slows continuous integration (CI) pipelines. In classical software, extensive research has shown that flakiness arises from diverse sources such as concurrency, nondeterministic execution order, platform dependencies, and environment-related timing issues (Gruber and Fraser, 2022; Parry et al., 2021). To mitigate these challenges, the classical literature has developed a spectrum of techniques: dynamic re-execution strategies (Alshammari et al., 2021; Lam et al., 2019b; Bell et al., 2018), static analysis (Luo et al., 2014; Gruber et al., 2021; Lam et al., 2020, 2019a; Parry et al., 2021), and machine-learning-based predictors (Ziftci and Cavalcanti, 2020; Bell et al., 2018; Lam et al., 2019b; Dutta et al., 2020; Alshammari et al., 2021; Verdecchia et al., 2021).
As the quantum computing ecosystem matures, similar reliability challenges have begun to surface. Frameworks such as Qiskit (Javadi-Abhari et al., 2024) now provide full software stacks and have testing infrastructures comparable in scale to classical systems. Yet, quantum software introduces additional sources of nondeterminism: probabilistic program semantics, noise-aware compilation flows, floating-point approximations, evolving toolchains, and constraint–sensitive optimizations in the transpiler. Early studies have identified flaky tests in quantum software repositories (Zhang et al., 2023), but current work relies primarily on static analysis small, manually curated incident-report-driven datasets (Zhang and Miranskyy, 2024; Kaur et al., 2025). Thus, we still lack a systematic, empirical understanding of how frequently flaky tests occur in quantum software, how they evolve across releases, and how difficult they are to detect given realistic rerun budgets.
Static machine-learning approaches depends critically on the availability of reliable ground truth. Without large, empirically derived datasets, especially those capturing very low-probability events, such models risk overfitting to limited historical cases (Verdecchia et al., 2021). Moreover, given that failure probabilities in quantum software can be extremely small, the limited rerun budgets typically used in CI pipelines may be insufficient to reveal rare nondeterministic behavior.
To close this gap, we conduct the first large-scale dynamic study of flaky tests in quantum software. We repeatedly execute the complete test suite for the Qiskit terra component [^3] From hereon we will refer to it as terra. Its role is summarized in the project documentation: “This library is the core component of Qiskit, which contains the building blocks for creating and working with quantum circuits, quantum operators, and primitive functions (Sampler and Estimator). It also contains a transpiler that supports optimizing quantum circuits, and a quantum information toolbox for creating advanced operators.” (Qiskit Development Community, 2025j) 10,000 times across 23 releases in controlled, containerized environments. This design surfaces rare nondeterministic failures that would remain undetected under conventional testing strategies and enables statistical analysis of empirical failure probabilities, temporal persistence, and subcomponent-level patterns.
We formally state the research questions (RQs) guiding this study.
  1. RQ1:
    How do the count and fraction of flaky tests change between terra releases?
  2. RQ2:
    How persistent are individual flaky test cases across terra releases?
  3. RQ3:
    What is the probability of detecting a flaky test given a specific number of executions?
  4. RQ4:
    How are failure intensities distributed across releases?
  5. RQ5:
    How are flaky tests distributed across terra subcomponents?
  6. RQ6:
    How do flaky tests emerge, dissipate, or recur across the sequence of terra releases?
This work makes the following contributions.
  1. 1.
    Large-scale empirical analysis. We perform the first dynamic investigation that reruns terra test suite 10,000 times across 23 terra releases (approximately 70 CPU-years of computation). This reveals how the prevalence and characteristics of flaky tests evolve over time.
  2. 2.
    Probabilistic characterization of detectability. We quantify empirical failure probabilities for flaky-tests and derive confidence-based estimates of the number of repetitions required to detect failures of varying rarity. Our analysis shows that many quantum flaky tests exhibit extremely sparse failures, requiring large number of reruns to surface reliably.
  3. 3.
    Public dataset of flaky tests. We release a curated dataset containing 290 unique quantum flaky tests, annotated with results from 10,000 controlled executions per release. This dataset provides high-quality ground truth for future research in quantum test reliability, debugging, and machine-learning prediction. It is available at https://zenodo.org/records/17979349.
By integrating large-scale dynamic evidence with probabilistic analysis, this study offers the first comprehensive view of flaky-test behavior in quantum software. The findings highlight the practical challenges in detecting low-probability nondeterminism and establish a foundation for more reliable quantum testing pipelines.
The remainder of the paper is organized as follows. Section 2 reviews related work. Section 3 describes the software under study, the data collection process, and the execution environment. Section 4 presents the empirical analysis and answers the RQs. Section 5 concludes.
2 Related Work
Flaky tests have been widely extensively in classical software, and only recently examined in quantum software. This section reviews both bodies of work, highlighting the methodological gap that motivates our dynamic study.
2.1 Flaky Test in Classical Software
Research on classical flaky tests spans three main areas: dynamic detection, static prediction, and root-cause analysis as discussed below.
Early work on classical flaky tests focused on detecting them by re-running tests under varying conditions or orders. Techniques like iDFlakies (Lam et al., 2019b) repeatedly reorder and execute test suites to surface order-dependent flaky behavior, resulting in a dataset of 422 flaky tests. DeFlaker (Bell et al., 2018) detects flaky failures in real-time and flags failures not associated with recent code changes. Such methods remain effective but are computationally expensive, especially for large test suites.
To reduce rerun cost, several tools predict flakiness from code characteristics. For instance, Pinto et al. (2020) present a purely static approach that warns developers about flaky tests as they write code. FLAST (Verdecchia et al., 2021) represents test code in a high-dimensional space using bag-of-words features and applies sparse random projection with $k$-nearest neighbors. FlakeFlagger (Alshammari et al., 2021) combines static and dynamic features, supported by a dataset created by running 811 tests 10,000 times — far larger than previous dynamic efforts. They show that small rerun budgets (10, 100, and 1000 re-runs) detect only a minority of flaky tests ($\approx$ 26%, 45%, and 67%, respectively), motivating higher-budget studies.
Luo et al. (2014) provide a widely used taxonomy of ten flaky-test root causes, with asynchronous waits, concurrency, and order dependencies accounting for most cases. Eck et al. (2019) discovered four new causes of flakiness, including those related to test environments and unmet assumptions that had not been reported previously.
2.2 Flaky test in Quantum Software
Flaky tests in quantum software have received limited attention, with existing research relying almost exclusively on static or manual methods.
Zhang et al. (2023); Zhang and Miranskyy (2024) performed the first empirical study of flaky tests in quantum repositories by mining issue trackers and pull requests across 14 open-source projects. They identified 46 instances of flaky tests in 12 projects (0.3–1.8% of reported bugs), categorized eight common causes of quantum test flakiness, and documented seven common fix strategies. This work demonstrates that flakiness already exists in quantum ecosystems, but the scale is restricted by the availability of human-reported incidents.
Kaur et al. (2025) proposed the first machine-learning-based flaky-test detector for quantum software. Using a bag-of-words representation similar to classical static predictors (Verdecchia et al., 2021), they trained various classifiers that achieved strong performance on the limited set of available labeled cases. However, like classical static approaches, the accuracy of such models is fundamentally constrained by the lack of large, dynamic ground-truth datasets.
Sivaloganathan et al. (2024) proposed an automated framework for detecting flaky tests in quantum software by mining issue reports and pull requests using embedding-based similarity and large language models, expanding prior manually curated datasets. While their approach achieves strong performance for flakiness detection from textual artifacts, it remains dependent on developer-reported incidents rather than direct observation of execution-level nondeterminism.
2.3 Summary and Gap
Across both domains, dynamic rerun studies have proven essential for uncovering rare nondeterministic failures and for building reliable ground-truth datasets. Yet, no such large-scale dynamic analysis exists for quantum software. Prior quantum studies rely on incident reports or static features and therefore cannot: (i) measure true flakiness prevalence across versions, (ii) quantify empirical failure probabilities, or (iii) assess detectability under realistic rerun budgets.
This work addresses that gap by conducting the first large-scale dynamic flakiness study in a quantum software framework. Our dataset provides the empirical foundation needed for future statistical analysis and ML-based prediction.
3 Study Design
This study aims to measure and characterize flaky behavior in quantum software through large-scale dynamic re-execution. This section describes the software under study, our data-collection pipeline, and the execution environment used to generate the dataset.
3.1 Software Under Study
We focus on the Qiskit software stack (Javadi-Abhari et al., 2024), one of the most widely used frameworks for quantum circuit construction, compilation, and analysis (Khan et al., 2025). Qiskit provides high-level circuit APIs, a multi-stage transpiler, simulators, backend interfaces, and a comprehensive built-in test suite, making it a representative and mature target for studying nondeterministic behavior in quantum software.
Our analysis concentrates on the core Qiskit library (Javadi-Abhari et al., 2024), formerly known as qiskit-terra prior to v.1.0, renamed qiskit thereafter. Its role is summarized in Footnote 3.
We study all 23 releases between v.0.25.0 and v.1.2.4 (see Table 2). The lower bound reflects an engineering constraint (the earliest versions rely on outdated toolchains that no longer build reproducibly) while the upper bound coincides with the last release supporting Python 3.8, which ensures compatibility across all versions under study.
3.2 Data Collection Method and Execution Testbed
Our goal is to obtain high-fidelity, release-specific measurements of nondeterministic test outcomes. To achieve this, we designed a pipeline that repeatedly executes terra test suite 10,000 times per release in clean, isolated environments.
Qiskit uses TOX software (Tox Development Community, ) to virtualize the test environment and execute test cases on specific platforms. Since Linux is both Qiskit’s primary development platform and the operating system of the high-performance computing (HPC) cluster that we used to run test cases, we selected it for our study.
Each run of the test suite was executed in a clean and isolated environment to ensure that the environment’s state did not affect the tests. For each version, we prepared a dedicated Dockerfile that pins Python 3.8 and installs all dependencies needed to build and run that version’s test suite. Because Qiskit depends on Rust toolchain (Matsakis and Klock, 2014), we selected a compatible Rust toolchain for each release based on Qiskit’s evolving minimum requirements. Although required libraries were cached in the Docker images, internet access was still necessary because tox, Qiskit’s test orchestrator, resolves and verifies dependencies at runtime before executing the test suites.
Our HPC environment supports Singularity (Kurtzer et al., 2017) rather than Docker containers. Accordingly, each Docker image was converted to a Singularity image using the Singularity CLI. We then scheduled 11,000 executions per image, each using 4 CPU cores (Intel Broadwell processors) and 24 GB of RAM. Gathering the data on our testbed required approximately 70 CPU-years.
The slight over-provisioning (11,000 runs) provides resilience against transient issues such as network failures or dependency-resolution timeouts. After execution, we retained only the first 10,000 successful runs per release to maintain strict comparability across versions.
Each test suite produces text-formatted output in the standard pytest style. We collected all outputs into version-specific folders and parsed them using custom Python scripts. These scripts extract per-test pass/fail/skip counts and consolidate them into a structured CSV file. This process yields one CSV table per release, each summarizing approximately 20,000 test cases.
3.3 Dataset Description
All collected artifacts are publicly available at https://zenodo.org/records/17979349. Each of the 23 CSV files (one per release) shares a common schema summarized in Table 1. The fields describe per-test execution outcomes and the environment used for that release.
Table 1: Schema of the per-release execution dataset. Each row corresponds to one test case in one terra release.
Attribute Description
test_case_name The identifier of the individual terra test case.
passed_attempts Number of runs in which the test passed.
failed_attempts Number of runs in which the test failed.
skipped_runs Number of skipped runs (e.g., via skip markers or environment constraints); skips were outside our control.
total_runs The total number of executions attempted for the test case; computed as $\texttt{passed\_attempts}+\texttt{failed\_attempts}+\texttt{skipped\_runs}$; equals 10,000 in all retained rows.
success_rate The fraction of successful executions; computed as $\texttt{passed\_attempts}/\texttt{total\_runs}$.
qiskit_version The terra release tag used for the run.
python_version Python version used (always 3.8).
rust_version Rust toolchain version used to satisfy terra build requirements for that release.
We annotated flakiness per terra version by analyzing the 10,000 execution results of each test. We label a test as flaky within a release if it exhibits both at least one pass and at least one fail: formally, $\texttt{passed\_attempts}>0$ and $\texttt{failed\_attempts}>0$. We then counted flaky tests per version as the number of test cases meeting this criterion.
4 Study Results
This section presents the empirical findings of our study. We analyze flakiness across the terra releases and address each research question, followed by a discussion of the threats to the validity of our findings.
Initial Observations
Across the 23 releases, we observed 820 flaky-test occurrences, corresponding to 290 unique flaky tests out of 27,026 distinct test cases. Table 2 summarizes the flakiness count, flakiness rate, and the relative contribution of each release to the global set of unique flaky tests.
The number of flaky tests varies substantially across versions. For instance, v.0.25.0 contains 88 flaky tests, the largest among all releases, while v.1.2.3 contains none. Flakiness percentages remain low compared with classical systems (0–0.40%).
Let us now focus on individual research questions.
Table 2: Test suite size and flakiness across terra releases. For each release, we report the total number of test cases, the number of non-skipped tests, the number of flaky tests (i.e., those that both pass and fail across 10,000 runs), the flakiness rate, and the contribution of that release to the global set of unique flaky tests. The flakiness rate is computed as the ratio of flaky test count over total test count. The share of all unique flaky tests is computed as the ratio of the flaky test count for a given release over the total number of unique flaky tests across all 290 releases. The final two rows aggregate totals across all releases.
Version Total Test Count Non-Skipped Tests Count Flaky Tests Count Flakiness rate (%) Share of Unique Flaky Tests (%)
0.25.0 $21,789$ $21,445$ 88 0.40 30.34
0.25.1 $21,796$ $21,303$ 74 0.34 25.51
0.25.2 $21,813$ $21,320$ 13 0.06 4.48
0.25.3 $21,834$ $21,339$ 14 0.06 4.83
0.45.0 $21,931$ $21,447$ 26 0.12 8.97
0.45.1 $21,938$ $21,455$ 34 0.15 11.72
0.45.2 $21,947$ $21,464$ 40 0.18 13.79
0.45.3 $21,947$ $21,464$ 38 0.17 13.10
0.46.0 $21,975$ $21,491$ 41 0.19 14.14
0.46.1 $21,979$ $21,494$ 42 0.19 14.48
0.46.2 $21,989$ $21,504$ 39 0.18 13.45
0.46.3 $21,994$ $21,509$ 22 0.10 7.59
1.0.0 $16,316$ $16,122$ 38 0.23 13.10
1.0.1 $16,320$ $16,126$ 36 0.22 12.41
1.0.2 $16,333$ $16,139$ 45 0.28 15.52
1.1.0 $17,406$ $17,208$ 30 0.17 10.34
1.1.1 $17,418$ $17,220$ 31 0.18 10.69
1.1.2 $17,428$ $17,229$ 25 0.14 8.62
1.2.0 $17,668$ $17,468$ 26 0.15 8.97
1.2.1 $17,678$ $17,474$ 70 0.40 24.14
1.2.2 $17,684$ $17,480$ 26 0.15 8.97
1.2.3 $17,684$ $17,480$ 0 0.00 0.00
1.2.4 $17,684$ $17,480$ 22 0.12 7.59
Total $452,551$ $444,661$ 820 0.18 100.0
Unique $27,026$ $26,639$ 290
4.1 RQ1: How do the count and fraction of flaky tests change between terra releases?
Table 2 summarizes the number of flaky tests per release across the 23 terra versions studied. We observe substantial variation: the number of flaky tests ranges from 0 (in v.1.2.3) to 88 (in v.0.25.0), with an average of 35.3 (SD = 19.7). Flakiness percentages span 0–0.40%, with a mean of 0.18% (SD = 0.11%).
Flakiness is not monotonic over time and shows episodic spikes. Early releases (v.0.25.0 and v.0.25.1) exhibit the highest absolute counts, while post-1.0 releases fluctuate in a narrower range but still show occasional spikes (e.g., v.1.2.1). Major-version transitions sometimes coincide with increases in flakiness (e.g., v.0.25.3 $\rightarrow$ v.0.45.0 or v.0.46.3 $\rightarrow$ v.1.0.0).
The complete absence of flakiness in v.1.2.3 is rare and is likely attributable to temporary stabilization rather than a systematic trend. Overall, the quantum test flakiness is episodic: it rises around major architectural or dependency changes, stabilizes for several releases, and occasionally resurfaces.
4.2 RQ2: How persistent are individual flaky test cases across terra releases?
The answer to the question is given in Table 3. We identified 290 unique flaky tests across the 23 releases. Most are short-lived and do not reappear: 192 tests (66.2%) appear in exactly one release. The remaining 98 tests (33.8%) recur across multiple releases. A small number of tests exhibit long recurrence spans, including: 8 (2.8%) tests flaky in 15 releases, 6 (2.1%) tests — in 14 releases, and 6 (2.1%) tests — in 12 releases. We will explore patterns of reappearance in Section 4.6.
Table 3: Persistence of flaky tests across terra releases. Each row shows the number of tests that were flaky in exactly $k$ terra releases within our study window (v.0.25.0–v.1.2.4). The maximum observed persistence was $k=15$, i.e., no test was flaky in more than 1515 releases.
Number of Releases ($k$) Flaky Tests Count Flaky Tests (%)
1 192 66.21
2 38 13.10
3 4 1.38
4 2 0.69
5 1 0.34
6 6 2.07
7 12 4.14
8 10 3.45
9 2 0.69
11 3 1.03
12 6 2.07
14 6 2.07
15 8 2.76
Total 290 100.00
4.3 RQ3: What is the probability of detecting a flaky test given a specific number of executions?
4.3.1 Empirical analysis
For each test case, we compute the empirical failure probability
$\hat{p}_{i}=\frac{F_{i}}{N},\qquad i=1,\ldots,290,$
where $F_{i}$ is the number of failed runs of test $i$ and $N=10000$ is the total number of runs of test $i$. Figure 1 shows the distribution of these estimates. Most (192/290) tests have $\hat{p}=1/10000=10^{-4}$, but the distribution exhibits a fat tail (median $\hat{p}=0.0033$), with some tests having $\hat{p}=9999/10000=0.9999$.
It is easy to miss test flakiness when $\hat{p}$ is low (e.g. $\hat{p}=10^{-4}$) under small rerun budgets. Let us perform a theoretical analysis to quantify the hardness.
Figure 1: Distribution of empirical failure probabilities $\hat{p}$ for all flaky tests in our dataset (290 unique tests, aggregated across releases). The distribution is fat-tailed: most tests concentrate at very small probabilities ($\hat{p}\approx 10^{-4}$), but there are a few that go to $\hat{p}=0.9999$.
4.3.2 Theoretical analysis
How many repetitions are needed to confidently detect a flaky test? Assuming independent and identically distributed test outcomes, failures follow a binomial distribution. To bound the true failure probability $p$ given an empirical estimate $\hat{p}$ at a chosen confidence level, we use the Wilson confidence interval for binomial proportions (Wilson, 1927); see Brown et al. (2001) for a review.
Figure 2: Wilson confidence intervals for the true failure probability $p$ as a function of the empirical estimate $\hat{p}$ when $N=10000$. For each confidence level $\gamma\in\{90\%,95\%,99\%\}$, the upper and lower bounds are shown as paired lines with matching color and line type. The reference line $p=\hat{p}$ is shown with a dot-dash pattern.
Figure 2 shows how the Wilson interval varies with $\hat{p}$ at different confidence levels $\gamma$. As $\hat{p}$ decreases, the interval widens.
We next connect these bounds to a conservative rerun budget for flakiness detection. Let $X$ be the number of observed failures after $n$ executions, so that $X\sim\mathrm{Binomial}(n,p)$. In general, the probability of observing exactly $k$ failures is
$\Pr(X=k)=\binom{n}{k}p^{k}(1-p)^{n-k},\qquad k=0,1,\dots,n.$
Therefore, the probability of detecting flakiness by observing at least one failure is
$\Pr(X\geq 1)=1-\Pr(X=0)=1-\binom{n}{0}p^{0}(1-p)^{n}=1-(1-p)^{n}.$ (1)
Given a target detection probability $q\in(0,1)$, we require
$1-(1-p)^{n}\geq q\quad\Leftrightarrow\quad(1-p)^{n}\leq 1-q\quad\Rightarrow\quad n\geq\frac{\ln(1-q)}{\ln(1-p)},$ (2)
as $\ln(1-p)<0$.
In practice, $p$ is unknown. Let $[L,U]$ be the Wilson confidence interval for $p$ with nominal coverage $\gamma$, computed from $N$ runs, and let $L=L_{\gamma}(\hat{p},N)$ denote its lower bound. Using $L$ as a conservative lower bound on the unknown $p$, and noting that Equation 1 is monotonically increasing in $p$, the worst case occurs at $p=L$. Therefore, a conservative rerun budget that targets detection probability at least $q$ is
$n_{q,\gamma}(\hat{p},N)=\left\lceil\frac{\ln(1-q)}{\ln\left[1-L_{\gamma}(\hat{p},N)\right]}\right\rceil,$
where $\lceil\cdot\rceil$ denotes the ceiling function.
Figure 3 plots $n_{q,\gamma}(\hat{p},N)$ versus $\hat{p}$ (for fixed $N$ and $q$), illustrating how required repetitions grow rapidly as $\hat{p}$ decreases and increase further with higher confidence $\gamma$. For example, with $q=0.95$ and $\gamma=0.95$, when $N=10000$ and $\hat{p}=10^{-4}$, then the 95% Wilson lower bound is $L_{0.95}\approx 1.77\times 10^{-5}$. Plugging these values into Equation 2, yields $n_{0.95,0.95}(0.0001,10000)\approx 1.69\times 10^{5}$ repetitions. When $\hat{p}=0.01$ and $L\approx 8.23\times 10^{-3}$, this reduces to $n_{0.95,0.95}(0.01,10000)\approx 363$ repetitions.
Figure 3: Conservative rerun budget for flakiness detection using Wilson lower confidence bounds. The $x$-axis shows the empirical estimate $\hat{p}$ from $N=10000$ observed runs, and the $y$-axis shows the required number of repetitions $n_{q,\gamma}(\hat{p},N)$ with $q=95\%$ and $\gamma\in\{90\%,95\%,99\%\}$. Higher confidence levels require more repetitions.
4.4 RQ4: How are failure intensities distributed across releases?
To characterize failure intensity within a release, we grouped empirical failure counts (out of 10,000 runs) into four bins: $(0,10]$, $(10,100]$, $(100,1000]$, and $(1000,10000]$. As shown in Figure 4, the distribution is dominated by the lowest bin, where $\hat{p}\leq 10^{-3}$. High-frequency flaky tests ($\hat{p}>0.1$) are rare but present. Notably, tests in the $(100,1000]$ bin appear relatively late, starting with v.0.45.1.
Figure 4: Distribution of failure frequencies for flaky tests across terra releases. Each horizontal bar partitions a release’s flaky tests into four bins based on the number of failures observed in 10,000 executions.
4.5 RQ5: How are flaky tests distributed across terra subcomponents?
To understand which parts of terra are most affected by nondeterminism, we mapped each test to its subcomponent. Subcomponent descriptions are provided in A. Figures 6 and 5 show that flakiness is highly nonuniform across the subcomponents.
Figure 5 indicates that the transpiler subcomponent has the highest number of distinct flaky tests (88), followed by quantum_info (66) and compiler (32). In contrast, subcomponents such as optflow and primitives each contain only 1 flaky tests.
Figure 6 provides a per-release breakdown, highlighting further variation across subcomponents. For example, the transpiler shows a median of 3 distinct flaky tests per release, with a maximum of 36. The qpy subcomponent, while present in 19 releases, has a narrower range, with a maximum of 12 and a median of 5 flaky tests per release.
Overall, these results suggest that the flakiness in terra is concentrated in a few key subcomponents, such as transpiler and quantum_info. The wide variation in flaky test frequency across releases also indicates that some components are more sensitive to changes in the codebase or external dependencies. Understanding this distribution may help prioritize stabilization efforts and guide targeted improvements in test reliability.
Figure 5: Total number of distinct flaky tests by terra subcomponent, aggregated across all releases. Subcomponents are sorted in descending order.
Figure 6: Distribution of flaky tests across terra subcomponents over all analyzed releases. For each subcomponent, the boxplot summarizes the per-release count of distinct flaky test cases, and subcomponents are ordered as in Figure 5, where $n$ indicates the number of releases in which flaky tests are observed for that subcomponent.
4.6 RQ6: How do flaky tests emerge, dissipate, or recur across the sequence of terra releases?
We analyzed time-series patterns of flaky appearances across releases and identified three characteristic temporal profiles.
Rarely Flaky
Tests flaky in $\leq 15\%$ of their active releases. This subtype dominates the dataset with 218 tests (75.2%). These failures are rare; examples are presented in Figure 7.
Persistently Flaky
Tests flaky in $\geq 70\%$ of the releases in which they appear. Only 18 tests (6.2%) fall into this group. They frequently fail but with variable intensity, often associated with unstable transformations in the transpiler. They often fail with fluctuating intensity; examples are presented in Figure 8.
Intermittently Flaky
Tests that alternate between flaky and stable states across multiple releases. This subtype includes 54 tests (18.6%) and is characterized by an intermittent, back-and-forth pattern where each flaky period is brief. Examples are presented in Figure 9.
The longest-streak distribution in Figure 10 shows that consecutive flakiness is uncommon: most tests never remain flaky across adjacent releases ($k=1$), indicating that recurrence is typically fragmented rather than sustained.
Figure 7: Rarely Flaky pattern. The temporal pattern in the right panel does not imply a sequence of flaky–fixed–reintroduced states. Rather, the underlying failure event is extremely rare, so the true failure probability is low and may not be consistently observed (see Section 4.3.2). Because a log scale cannot display zero, we use $10^{-6}$ as a placeholder for zero observed failures. The lowest observable non-zero rate is $10^{-4}$, corresponding to one failure in 10,000 runs.
Figure 8: Persistently Flaky pattern. The gray background marks releases in which the test case did not exist or was skipped. In the left panel, the test is present across several releases and consistently flaky until removal. In the right panel, the test is introduced later, remains consistently flaky, and appears to be fixed only in the final two releases. Similar to Figure 7, as log scales cannot display zero, we use $10^{-6}$ for zero observed failures; the lowest observable non-zero rate is $10^{-4}$.
Figure 9: Intermittently Flaky pattern. These examples show test cases that intermittently alternate between flaky and stable states. We do not speculate on the underlying causes, but this pattern contrasts with consistently failing tests such as those in Figure 8. As in Figures 7 and 8, $10^{-6}$ denotes zero observed failures, and the smallest observable non-zero rate is $10^{-4}$.
Figure 10: Longest consecutive flaky-release streak per test across terra releases. Bars show the number of flaky tests whose maximum streak length is $m$; most tests have $m=1$, while long streaks ($m\geq 2$) are rare.
4.7 Threats to Validity
We discuss threats to the validity of our case study following the classifications of Wohlin et al. (2012); Yin (2009).
4.7.1 Internal
Our methodology is derived from the classical experimental methodology of Flake Flagger (Alshammari et al., 2021): a high-budget rerun protocol (10,000 executions per test per release) in a controlled environment without injecting additional non-determinism. While this approach provides robust empirical data, it also introduces several risks.
First, repeated executions on an HPC cluster may not be independent or identically distributed due to factors like CPU frequency scaling, file system cache warm-up, and job contention, which can cause empirical failure rates to drift within a batch.
Second, unlike Lam et al. (2019b); Bell et al. (2018), we do not reorder tests, perturb timing, or inject failures. As a result, we may under-detect order-dependent and timing-sensitive flaky tests.
Although we pin Python (3.8) and specific versions of the Qiskit and Rust toolchains, some transitive dependencies are resolved at runtime. Changes in mirror or point-release updates can introduce subtle environment differences over time. To mitigate this, we containerize each version and fix the test orchestrator configuration. However, full immutability is difficult to guarantee without fully offline and locked dependencies.
4.7.2 External
Our results are based on the terra component running on Linux within Singularity containers on Intel-based HPC nodes. Hardware characteristics (e.g., ISA, cache hierarchy), operating system behavior, container runtime, and CI topology all influence timing and scheduling. Other environments (e.g., macOS/Windows, AMD CPUs, cloud-based CI) may exhibit different flakiness patters.
We restricted the release range to v.0.25.0–v.1.2.4 due to toolchain constraints. Earlier or later versions, or other quantum computing frameworks, may behave differently (Zhang et al., 2023). Therefore, generalization beyond Qiskit, beyond Python, or to hardware backends (as opposed to simulators) should be approached with caution.
As in many software engineering studies, the variability of real-world environments limits generalizability (Wieringa and Daneva, 2015). Nonetheless, our empirical methodology is reproducible and can be applied to other frameworks. We encourage the community to replicate and extend our work.
5 Conclusions and Future Work
This paper presents the first dynamic study of flaky tests in quantum software. By re-executing the Qiskit terra test suite 10,000 times across 23 releases, we provide a comprehensive empirical characterization of nondeterministic behavior in a quantum software testing framework.
Our findings show that test flakiness in quantum software is rare but often episodic. Although overall flakiness rates (0–0.40%) are lower than those typically reported in classical systems, we identified 290 unique flaky tests, with roughly two-thirds appearing in only a single release. At the same time, a nontrivial minority exhibit long or intermittent recurrence patterns, including persistently flaky tests that fail in most releases and intermittently flaky tests that alternate between stable and unstable states. This diversity indicates that quantum flakiness, like its classical counterpart, manifests through multiple behavioral modes.
Our probabilistic analysis further demonstrates that many quantum flaky tests fail with extremely small empirical probabilities ($\hat{p}=10^{-4}$). Detecting such infrequent failures with statistical confidence requires tens of thousands of executions, well beyond the rerun budgets typically available in CI pipelines. This highlights a structural challenge for quantum software testing: many flaky tests may evade detection unless substantially larger or adaptive rerun budgets are used.
We also release a publicly available dataset containing per-test execution results for all 23 releases, providing dynamic ground truth for future work in quantum software testing, debugging, and machine-learning prediction.
Future work includes (i) extending our analysis to other quantum software frameworks, (ii) exploring more efficient or adaptive detection strategies, and (iii) developing machine-learning models that leverage both static and dynamic signals for more robust detection and prediction of quantum flaky tests.
Acknowledgement
This work was partially supported by the Natural Sciences and Engineering Research Council of Canada (grant # RGPIN-2022-03886). The authors thank the Digital Research Alliance of Canada for providing computational resources.
Appendix A Module taxonomy
Below are descriptions of the main subcomponents of Qiskit terra.
transpiler
Tests passes, preset pass managers, layout, routing, basis translation, scheduling, and transpile() behavior. This subcomponent is present and actively maintained across all studied releases (0.25.0–1.2.4) (Qiskit Development Community, 2025p).
quantum_info
Tests the state/operator/channel toolbox (e.g., Statevector, DensityMatrix, SparsePauliOp), including conversions, algebraic operations, metrics, and channel properties. Available throughout 0.25.0–1.2.4 (Qiskit Development Community, 2025e, o).
compiler
Tests high-level compile wrappers (e.g., transpile, assemble) that orchestrate transpilation, scheduling, and backend configuration. Present as a thin wrapper interface in all considered versions (Qiskit Development Community, 2025d).
synthesis
Tests unitary and operator synthesis routines that decompose targets into supported gate sets (exact and approximate). Supported across 0.25.0–1.2.4 (Qiskit Development Community, 2025m).
circuit
Tests the core QuantumCircuit abstraction, gate and instruction definitions, parameterization, control-flow, and circuit transformation APIs. Central and stable over 0.25.0–1.2.4 (Qiskit Development Community, 2025c).
qpy
Tests QPY-based serialization and deserialization of circuits, checking round-trip correctness and compatibility. Used across all studied releases (Qiskit Development Community, 2025i).
pulse
Tests pulse-level programming (channels, waveforms, schedule blocks) and alignment with backend timing/constraints. Available throughout 0.25.0–1.2.4, though some features are later streamlined (Qiskit Development Community, 2025h).
scheduler
Tests legacy helpers converting circuits to scheduled pulse programs, including timing and resource constraints. Relevant only to earlier releases in our range (removed around 0.45) (Qiskit Development Community, 2025l).
visualization
Tests circuit drawers, state plots, device/heatmap plots, and related options. Present from 0.25.0 to 1.2.4 (Qiskit Development Community, 2025q).
providers
Tests provider and backend interfaces (including fake backends), job submission, configuration, and result formats. Present across all studied versions (Qiskit Development Community, 2025g).
test_user_config
Tests loading and applying user configuration (e.g., defaults, visualization settings) from local config files. Present in our full range (Qiskit Development Community, 2025a).
basicaer
Tests legacy local simulators for correctness and compatibility. Only applicable to the subset 0.25.0–0.46.x within our window (Qiskit Development Community, 2025b).
opflow
Tests operator-flow abstractions for operator expressions, expectations, and gradients (now legacy). Appears in 0.25.0–0.46.x within our range (Qiskit Development Community, 2025n).
primitives
Tests the Sampler and Estimator primitives, including interfaces, result formats, and backend/runtime integration. Becomes relevant from roughly 0.39.0 and remains central through 1.2.4 (Qiskit Development Community, 2025f).
test_examples
Tests example scripts and tutorials to ensure that published examples run end-to-end against the public API. Present throughout 0.25.0–1.2.4 (Qiskit Development Community, 2025k).